OpenStudy (anonymous):

Darlene bought some 30 cent and some 60 cent candy bars. Altogether she bought 10 candy bars and paid a total of $4.20. How many candy bars at each price did she buy?

OpenStudy (anonymous):

.3x+.6y = 4.2 x+y = 10 x = 10-y 3-.3y+.6y=4.2 .3y = 1.20 y = 4 candy bars x = 10-y x = 10 - 4 = 6 candy bars